Mrs. De Guzman is 28 years old when her daughter was born. In how many years will Mrs. De Guzman be thrice as old as her daughter.
Found 2 solutions by josgarithmetic, ikleyn:
Answer by josgarithmetic(39620)
Answer by ikleyn(52803)

.

Let x be "in how many years . . . ".


Then you have this equation


    x + 28 = 3x.


The left side is the mother's age in 28 years; the right side is her age then, expressed via the daughter's age.


From the equation


    28 = 3x - x

    28 = 2x

    x  = 28/2 = 14.


ANSWER.  In 14 years.


CHECK.  In 14 years, the mother will be 28 + 14 = 42 years old;

                     the daughter will be         14 years old.   ! Ok, good, i.e. CORRECT !

Solved.

--------------

There is a bunch of lessons on age word problems
    - Age problems and their solutions
    - HOW TO algebreze and to solve age problems?
    - A fresh formulation of a traditional age problem
    - Really intricate age word problems
    - Selected age word problems from the archive
    - Age problems for mental solution
    - Age problem for three participants
    - Miscellaneous age problems
in this site.

Read them and become an expert in solving age problems.

Also,  you have this free of charge online textbook in ALGEBRA-I in this site
    - ALGEBRA-I - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Age word problems".


Save the link to this online textbook together with its description

Free of charge online textbook in ALGEBRA-I
https://www.algebra.com/algebra/homework/quadratic/lessons/ALGEBRA-I-YOUR-ONLINE-TEXTBOOK.lesson

to your archive and use it when it is needed.